Please help with these

1. Calculate ΔG (in kJ) at 298 K for some solid ZnF2, 0.082 M Zn2+ and 0.039 M F-(aq).

2.Calculate the equilibrium constant at 45 K for a reaction with ΔHrxno = 10 kJ and ΔSrxno = -100 J/K. (Don't round unil the end. Using the exponent enlarges any round-off error.

Calculate the equilibrium constant at 136 K for the thermodynamic data in the previous question. ( Problem 2)
Answer #1

1) ∆G = -RTlnK

K = [Zn 2+][F-]2 = (0.082)(0.039)2 = 1.25×10-4

T = 298K

R = 8.314×10-3 KJ/mol.K

∆G = -8.314×10-3×298×ln(1.25×10-4)

= 22.27 KJ

2) ∆G = ∆H - T∆S = 10 - 45(-100/1000) KJ

= 10 + 4.5 = 14.5 KJ

K = e​​​​​​-(∆G/RT) = e-(14.5/0.008314×45) = 1.47×10-17

3) ∆G = 10-136(-0.1) = 23.6 KJ

K = e-23.6/0.008314×136 = 8.64×10-10